# Modeling extremal dependence in multivariate and spatial problems: a practical perspective

### v3

@@ Abstract
From environmental sciences to finance, there is a growing demand for methods that can assess the risks of extreme events beyond those observed in available data. Extrapolating extreme events beyond the range of the data is not obvious and requires advanced tools based on extreme value theory. Risk assessments are often further complicated by the need to account for multiple variables simultaneously. Extreme value theory provides very important tools for the analysis of multivariate or spatial extreme events, but these are not easily accessible to professionals without appropriate expertise. This article provides a minimal background on multivariate and spatial extremes and gives simple yet thorough instructions on how to analyse them using the R package ExtremalDep. After briefly introducing the statistical methodologies, we focus on road testing the package’s toolbox through several real-world applications.
@@ 1 Introduction
In many applied fields—including environmental sciences, finance, and actuarial sciences—it is essential to assess the risk of extreme events beyond those observed in historical data, with the goal to anticipate the magnitude of hypothetical future catastrophes (e.g., a global financial crisis or large monetary losses due to natural hazards). The aim is to disclose the uncertainty surrounding such extreme events to decision makers in order to carefully plan mitigation strategies. A key challenge arises when extrapolating beyond the range of $N$ available observations, by computing for example quantiles whose probability ($p$) is smaller than $1/N$, meaning that the largest data observation is not expected to exceed the target event on average. This makes direct estimation impossible. Extreme Value Theory (EVT) addresses this problem by providing a rigorous probabilistic framework and statistical methodology for modeling and extrapolating extreme events. For accessible introductions, see, for instance, Beirlant et al. (2004), de Haan and Ferreira (2006), and Falk et al. (2010).
@@
In the univariate case, the theory and practice is mature with a significant number of packages available on the Comprehensive R Archive Network (CRAN), forming an extensive set of tools. Examples include ismev (Heffernan and Stephenson, 2018), evd Stephenson (2002), evdbayes (Stephenson and Ribatet, 2023), extRemes (Gilleland and Katz, 2016) and ExtremeRisks (Pacifici et al., 2025) to name a few. In practice, extreme events rarely involve a single variable. Multiple factors are typically observed simultaneously—for example, rainfall, temperature, and wind measured across different locations within a region can jointly drive economic losses from adverse weather. Similarly, high pollution levels often result from the combined effect of several pollutants exceeding safety thresholds.
@@
When it comes to multivariate and spatial problems, statistical methodologies for analyzing multivariate and spatial extremes—and their corresponding software implementations—remain relatively underdeveloped. Extremal dependence refers to the dependence structure governing multivariate extreme value distributions, or equivalently, the finite-dimensional distributions of stochastic processes for extremes. This dependence structure is an infinite-dimensional, nonparametric object subject to restrictive conditions, which leads to both a complex formulation of the overall multivariate distribution and significant inferential challenges. The literature on multivariate and spatial extremes covers several topics including multivariate extremes based on componentwise maxima (e.g., Beranger and Padoan, 2015) and multiple peaks over threshold (e.g., Rootzén et al., 2018), spatial extremes based on pointwise maxima and peaks over threshold (e.g., Davison et al., 2012; de Fondeville and Davison, 2018), graphical models (Engelke and Hitz, 2020) and time series (e.g., Davis et al., 2011), and remains nowadays a vibrant research field with a number of ongoing new results. These have produced useful R packages such as mev (Belzile et al., 2022), SpatialExtremes (Ribatet, 2022), mvPot (de Fondeville and Belzile, 2021), RandomFields, graphicalExtremes (Engelke et al., 2024), extremogram (Frolova and Cribben, 2016) as well as ExtremalDep (Beranger et al., 2025). A review of statistical software for modeling extremes is provided by Belzile et al. (2023), which introduces some of the capabilities of ExtremalDep; but we offer here a more in-depth guide to the package, complemented by practical illustrations.
@@
The focus of this work is on modeling the extremal dependence structure arising from the componentwise maxima approach, which enables the assessment of risk associated with multivariate extreme events in a broad sense. Contrary to what might be expected, this framework is not limited to modeling multivariate maxima, but is also well suited for analyzing multiple threshold exceedances. This is achieved by exploiting approximations of the probabilities of observing at least one exceedance (2.9), joint exceedances (2.10), or an observation falling within an extreme region (2.12), in which at least one component is extreme. Statistical methods for risk analysis based on these probability approximations are presented in Sections 2.3, 2.4.3, and 2.4.4. If the primary interest lies instead in proper multivariate peaks-over-threshold models and methods specifically tailored to their analysis, we refer the reader to Rootzén et al. (2018); Kiriliouk et al. (2019). Finally, it is worth noting that several alternative approaches for modeling multivariate extremes also exist, including conditional extremes models, random scale and location mixture models, and (X-)vine copulae (e.g., Heffernan and Tawn, 2004; Huser et al., 2017; Kiriliouk et al., 2025). Several of these alternative methods are implemented in the texmex package (Southworth et al., 2024).
@@
Most existing packages for the analysis of multivariate extremes rely on parametric representations of extremal dependence (e.g., mev, mvPot, and graphicalExtremes), which simplify the underlying mathematics and inference but impose strong structural assumptions. While ExtremalDep also accommodates parametric approaches, one of its main contributions lies in extending the available toolkit by providing nonparametric and semiparametric methods that remain closer to the inherently nonparametric nature of extremal dependence. A further distinguishing feature of ExtremalDep, compared with existing software, is its emphasis on demonstrating how extremal dependence can be directly exploited for risk assessment related to multivariate extreme events and their occurrence probabilities. In particular, the routines of the package allow to compute joint return levels and conditional return periods (Section 2.3), conditional probabilities given the occurrence of extreme events (Section 2.4.2), tail probabilities relative to multiple threshold exceedances (Section 2.4.3), and extreme quantile regions (Section 2.4.4). Finally, the package allows for frequentist and Bayesian inference for multivariate extremes and frequentist and Bayesian inference for spatial extremes. This manuscript demonstrates the practical use of the package through several important applications. Because the ordering of points in $d$-dimensional Euclidean space is arbitrary, we introduce and formalize key concepts such as extreme sets, multivariate extreme quantiles, quantile regions, joint tail and conditional probabilities, and multivariate return levels. In practice, these tools enable a wide range of risk analyses—for example, risk assessment of future high-pollution episodes in urban areas, identifying regions of a country exposed to heavy precipitation events, or evaluating the risk of future heatwave episodes. The article is divided into two parts, Section 2 discusses multivariate extremes while Section 3 focuses on spatial extremes. In the multivariate part, Section 2.1 gives a brief introduction to the theoretical background while Section 2.3 details how to perform parametric (frequentist and Bayesian) inference for the extremal dependence and for joint and conditional tail probabilities. Section 2.4 explains how to perform nonparametric Bayesian inference for the extremal dependence of a $d$-dimensional componentwise maxima vector, and once the extremal dependence has been estimated, how to infer joint and conditional tail probabilities in the bivariate case. The focus then successively shifts to the random generation of bivariate extreme values with a flexible semi-parametric dependence structure, the estimation of small probabilities of belonging to certain extreme sets and the estimation extreme quantile regions corresponding to small joint tail probabilities. The paper ends in Section 4 with a summary and a discussion on future developments.
@@ 2.1 Background
Let ${\bm{X}}_{1},{\bm{X}}_{2},\ldots$ be independent and identically distributed (i.i.d.) random vectors, where ${\bm{X}}_{i}=(X_{i,1},\ldots,X_{i,d})\in{\mathbb{R}}^{d}$ for some integer $d>1$, whose joint distribution $F$ is in the domain of attraction of a multivariate extreme value distribution $G$ (de Haan and Ferreira, 2006, chap. 6), shortly denoted as $F\in\mathcal{D}(G)$. This implies that for $n=1,2,\ldots$, there exists norming sequences ${\bm{a}}_{n}>{\bf 0}=(0,\ldots,0)$ where the inequality is applied element-wise, and ${\bm{b}}_{n}\in{\mathbb{R}}^{d}$ such that
@@
where $\gamma_{j}\in{\mathbb{R}}$ is the extreme value index. This coefficient describes the tail heaviness of the distribution, resulting in a heavy-, light- or short-tailed distribution, depending on whether $\gamma_{j}>0$, $\gamma_{j}=0$ or $\gamma_{j}<0$ (e.g., de Haan and Ferreira, 2006, chap. 1). Then, a crucial step in the computation of (2.2) is to know the form of extreme-value copula, which turns out to be defined as
@@
where $L:[0,\infty)^{d}\mapsto[0,\infty)$ is a homogeneous function of order $1$ named stable-tail dependence function (Beirlant et al., 2004, chap. 8). This latter function fully characterizes the extremal dependence structure; however, since it is defined on $0,\infty)^{d}$ and takes values in $0,\infty)$, it is difficult to visualize directly. For this reason, it is common practice to work instead with its restriction to the set ${\mathcal{R}}:=\{{\bm{t}}\in[0,1]^{d-1}:\lVert{\bm{t}}\rVert_{1}\leqslant 1\}$, known as the Pickands dependence function (e.g., Falk et al., 2010, Ch. 4), which is more amenable to graphical representation. A central question is the form of the Pickands dependence function. While no exhaustive characterization is available, it must at least satisfy the following convexity and boundary conditions
@@
In the bivariate case these are necessary and sufficient to define a valid Pickands dependence function (for details see Beirlant et al., 2004, chap. 8). However, there is an alternative way to represent extreme dependence. Indeed, a formal definition of the Pickands dependence function is
@@
where $H$ is a probability measure on the $d$-dimensional unit simplex $\mathcal{S}_{d}:=\{{\bm{w}}\geqslant{\bf 0}:\|{\bm{w}}\|_{1}=1\}$}, named angular measure (e.g., Falk et al., 2010, Ch. 4). Thus, the level of dependence can be conveyed through $H$; however, the same issue arises again: which specific measure should one consider? The answer is, a valid angular measure is any probability measure satisfying the mean constraint
@@
The result in (2.1) can be interpreted as follows. Given a sample $({\bm{X}}_{1},\ldots,{\bm{X}}_{N})$ with joint distribution $F\in\mathcal{D}(G)$, for each of the $d$ variables we compute $k$ maxima are computed over blocks of size $n$, such that $N=n\cdot k$, where the $i$th maximum is $M_{i,j}^{(n)}=\max(X_{(i-1)n+1,j},\ldots,X_{in,j})$, $i\in\{1,\ldots,k\}$, $j=1,\ldots,d$. We bind the maxima in the vector of componentwise maxima ${\bm{M}}_{i}^{(n)}=(M_{i,1}^{(n)},\ldots,M_{i,d}^{(n)})^{\top}$. Now, setting ${\bm{\mu}}={\bm{b}}_{n}$, ${\bm{\sigma}}={\bm{a}}_{n}$, ${\bm{y}}={\bm{a}}_{n}{\bm{x}}+{\bm{b}}_{n}$, then thanks to (2.1), the joint distribution of vector of componentwise maxima can be approximated, for large enough $n$, as $F^{n}({\bm{y}})\approx G_{{\bm{\theta}}}({\bm{y}}\mid D),$ where ${\bm{\theta}}=({\bm{\mu}},{\bm{\sigma}},{\bm{\gamma}})^{\top}$, By construction, the distribution of the univariate maximum $M_{i,j}^{(n)}$ can be approximated as
$F^{n}(y_{j})\approx G_{\theta_{j}}(y_{j})=\exp\left\{-\left(1+\gamma_{j}\frac{y_{j}-\mu_{j}}{\sigma_{j}}\right)^{-\frac{1}{\gamma_{j}}}_{+}\right\},\qquad j\in\{1,\ldots,d\}.$ (2.6)
@@
Recall that the joint distribution function $F$ can be represented in terms of a copula function, $C$, defined on $0,1]^{d}$ and its marginal distributions $F_{i},i=1,\ldots,d$ (assumed to be continuous) through $F({\bm{x}})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d}))$ (see e.g., Joe, 2014, for details). The limiting result in (2.8) implies the following important approximation: let $p_{j}=x_{j}/n$, for $j=1,\ldots,d$, and $Q_{j}(p_{j})$ be the (1p1-p) quantile of $F_{j}$; then, by homogeneity of the stable tail dependence function, the probability that at least one among the $d$ variables $X_{j}$ exceeds a high quantile of its own distribution, can be approximated for large $n$ by
${\mathbb{P}}(X_{1}>Q_{1}(p_{1})\,\text{or}\cdots\text{or}\,X_{d}>Q_{d}(p_{d}))\approx L(p_{1},\ldots,p_{d}).$ (2.9)
@@
${\mathbb{P}}(X_{1}>Q_{1}(p_{1})\,\text{and}\cdots\text{and}\,X_{d}>Q_{d}(p_{d}))\approx R(p_{1},\ldots,p_{d}).$ (2.10)
@@
Finally, a related question is, how to determine an extreme region, where two variables, given a small joint probability $p$ of falling in it? We summarize the approach discussed by Cai et al. (2011), Einmahl et al. (2013) and He and Einmahl (2017) as follows with a focus on the bivariate case explored by Einmahl et al. (2013). Let $f$ be the density on ${\mathbb{R}}_{+}^{2}$ of a bivariate distribution $F$ satisfying $F\in\mathcal{D}(G)$, with $\gamma_{1},\gamma_{2}>0$. A possible definition of quantile regions is given by the level sets of $f$, namely for $\alpha>0$
@@
Assume that: $f$ is bounded away from zero on $(0,\mathcal{M}]^{2}$ and, outside this region, $f$ is decreasing in each coordinate, for some $\mathcal{M}>0$; that the exponent measure admits a density $g$ satisfying
@@
where $k=k_{N}$ with $k\to\infty$ as $N\to\infty$ and $k=o(N)$, which approximate the target extreme quantile region $\mathcal{Q}_{N}$. In particular, $\mu_{j}=Q_{j}(k/N)$ and $\sigma_{j}=a_{j}(N/k)$ with $j=1,2$, where $a_{j}(\cdot)$ is a suitable positive scale function (de Haan and Ferreira, 2006, Chapters 1–2). $\widetilde{\mathcal{Q}}_{N}$ is a good approximation of $\mathcal{Q}_{N}$ in the sense that ${\mathbb{P}}(\widetilde{\mathcal{Q}}_{N})\approx p$ and ${\mathbb{P}}(\mathcal{Q}_{N}\Delta\widetilde{\mathcal{Q}}_{N})\to 0$ as $N\to\infty$ (see Einmahl et al., 2013), where $A\Delta B=A\setminus B\cup B\setminus A$.

### v5

@@ Abstract
From environmental sciences to finance, there is a growing demand for methods that can assess the risks of extreme events beyond those observed in available data. Extrapolating extreme events beyond the range of the data is not obvious. Risk assessments are often further complicated by the need to account for multiple variables simultaneously. Extreme value theory provides important tools for the analysis of multivariate or spatial extreme events, but these are not easily accessible to professionals without appropriate expertise. This article provides a minimal background on multivariate and spatial extremes and gives simple yet thorough instructions on how to analyse them using the R package ExtremalDep. After briefly introducing the statistical methodologies, we focus on road testing the package’s toolbox through several real-world applications.
@@ 1 Introduction
In many applied fields—including environmental sciences, finance, and actuarial sciences—it is essential to assess the risk of extreme events beyond those observed in historical data, with the goal to anticipate the magnitude of hypothetical future catastrophes (e.g., a global financial crisis or large monetary losses due to natural hazards). The aim is to disclose the uncertainty surrounding such extreme events to decision makers in order to carefully plan mitigation strategies. A key challenge arises when extrapolating beyond the range of $N$ available observations, by computing for example quantiles whose probability ($p$) is smaller than $1/N$, meaning that the largest data observation is not expected to exceed the target event on average. This makes direct empirical estimation impossible. Extreme Value Theory (EVT) addresses this problem by providing a rigorous probabilistic framework and statistical methodology for modeling and extrapolating extreme events. For accessible introductions, see, for instance, Beirlant et al. (2004), de Haan and Ferreira (2006), and Falk et al. (2010).
@@
In the univariate case, the theory and practice is mature with a significant number of packages available on the Comprehensive R Archive Network (CRAN), forming an extensive set of tools. Examples include ismev (Heffernan and Stephenson, 2018), evd Stephenson (2002), revdbayes (Northrop and Grimshaw, 2025), extRemes (Gilleland and Katz, 2016) and ExtremeRisks (Pacifici et al., 2025) to name a few. In practice, extreme events rarely involve a single variable. Multiple factors are typically observed simultaneously: for example, rainfall, temperature, and wind measured across different locations within a region can jointly drive economic losses from adverse weather. Similarly, high pollution levels often result from the combined effect of several pollutants exceeding safety thresholds.
@@
When it comes to multivariate and spatial problems, statistical methodologies for analyzing multivariate and spatial extremes, and their corresponding software implementations, remain relatively underdeveloped. The dependence structure governing multivariate extreme value distributions, or equivalently, the finite-dimensional distributions of stochastic processes for extremes is termed extremal dependence. It is an infinite-dimensional, nonparametric object subject to some restrictive conditions, which leads to both a complex formulation of the overall multivariate distribution and significant inferential challenges. The literature on multivariate and spatial extremes covers several topics including multivariate extremes based on componentwise maxima (e.g., Beranger and Padoan, 2015) and multiple peaks over threshold (e.g., Rootzén et al., 2018), spatial extremes based on pointwise maxima and peaks over threshold (e.g., Davison et al., 2012; de Fondeville and Davison, 2018), graphical models (Engelke and Hitz, 2020) and time series (e.g., Davis et al., 2011), and remains nowadays a vibrant research field with a number of ongoing new results. These have produced useful R packages such as mev (Belzile et al., 2025), SpatialExtremes (Ribatet, 2022), mvPot (de Fondeville et al., 2025), RandomFields, graphicalExtremes (Engelke et al., 2025), extremogram (Frolova and Cribben, 2016) as well as ExtremalDep (Beranger et al., 2025). A review of statistical software for modeling extremes is provided by Belzile et al. (2023), which introduces some of the capabilities of ExtremalDep; but we offer here a more in-depth guide to the package, complemented by practical illustrations.
@@
The focus of this work is on modeling the extremal dependence structure arising from the componentwise maxima approach, which enables the assessment of risk associated with multivariate extreme events in a broad sense. Contrary to what might be expected, this framework is not limited to modeling multivariate maxima, but is also well suited for analyzing multivariate threshold exceedances. This is achieved by exploiting approximations of the probabilities of observing at least one exceedance (2.9), joint exceedances (2.10), or an observation falling within an extreme region (2.12), in which at least one component is extreme. Statistical methods for risk analysis based on these probability approximations are presented in Sections 2.3, 2.4.3, and 2.4.4. If the primary interest lies instead in proper multivariate peaks-over-threshold models and methods specifically tailored to their analysis, we refer the reader to Rootzén et al. (2018); de Fondeville and Davison (2018); Kiriliouk et al. (2019). Finally, it is worth noting that several alternative approaches for modeling multivariate extremes also exist, including conditional extremes models (Heffernan and Tawn, 2004), random scale and location mixture models (Huser et al., 2017), and (X-)vine copulae (Kiriliouk et al., 2025). Several of these alternative methods are implemented in the texmex package (Southworth et al., 2024).
@@
Most existing packages for the analysis of multivariate extremes rely on parametric representations of extremal dependence (e.g., mev, mvPot, and graphicalExtremes), which simplify the underlying mathematics and inference but impose strong structural assumptions. While ExtremalDep also accommodates parametric approaches, one of its main contributions lies in extending the available toolkit by providing nonparametric and semiparametric methods that remain closer to the inherently nonparametric nature of extremal dependence. A further distinguishing feature of ExtremalDep, compared with existing software, is its emphasis on demonstrating how extremal dependence can be directly exploited for risk assessment related to multivariate extreme events and their occurrence probabilities. In particular, the routines of the package allow one to compute joint return levels and conditional return periods (Section 2.3), conditional probabilities given the occurrence of extreme events (Section 2.4.2), tail probabilities relative to multiple threshold exceedances (Section 2.4.3), and extreme quantile regions (Section 2.4.4). Finally, the package allows for frequentist and Bayesian inference for multivariate extremes and frequentist and Bayesian inference for spatial extremes. This manuscript demonstrates the practical use of the package through several important applications. Because the ordering of points in $d$-dimensional Euclidean space is arbitrary, we introduce and formalize key concepts such as extreme sets, multivariate extreme quantiles, quantile regions, joint tail and conditional probabilities, and multivariate return levels. In practice, these tools enable a wide range of risk analyses. Examples include risk assessment of future high-pollution episodes in urban areas, identifying regions of a country exposed to heavy precipitation events, or evaluating the risk of future heatwave episodes. The article is divided into two parts: Section 2 discusses multivariate extremes while Section 3 focuses on spatial extremes. In the multivariate part, Section 2.1 gives a brief introduction to the theoretical background while Section 2.3 details how to perform parametric (frequentist and Bayesian) inference for the extremal dependence and for joint and conditional tail probabilities. Section 2.4 explains how to perform nonparametric Bayesian inference for the extremal dependence of a $d$-dimensional componentwise maxima vector, and once the extremal dependence has been estimated, how to infer joint and conditional tail probabilities in the bivariate case. The focus then successively shifts to the random generation of bivariate extreme values with a flexible semi-parametric dependence structure, the estimation of small probabilities of belonging to certain extreme sets and the estimation extreme quantile regions corresponding to small joint tail probabilities. The paper ends in Section 4 with a summary and a discussion on future developments.
@@ 2.1 Background
Let ${\bm{X}}_{1},{\bm{X}}_{2},\ldots$ be independent and identically distributed (i.i.d.) random vectors, where ${\bm{X}}_{i}=(X_{i,1},\ldots,X_{i,d})\in{\mathbb{R}}^{d}$ for some integer $d>1$, whose joint distribution $F$ is in the domain of attraction of a multivariate extreme value distribution $G$ (de Haan and Ferreira, 2006, chap. 6), denoted by $F\in\mathcal{D}(G)$. This implies that for $n=1,2,\ldots$, there exists norming sequences ${\bm{a}}_{n}>{\bf 0}_{d}=(0,\ldots,0)$ where the inequality is applied element-wise, and ${\bm{b}}_{n}\in{\mathbb{R}}^{d}$ such that
@@
where $\gamma_{j}\in{\mathbb{R}}$ is the extreme value index. This coefficient describes the tail heaviness of the distribution, resulting in a heavy-, light- or short-tailed distribution, depending on whether $\gamma_{j}>0$, $\gamma_{j}=0$ or $\gamma_{j}<0$ (e.g., de Haan and Ferreira, 2006, chap. 1). In (2.2) the extreme-value copula is defined as
@@
where $L:[0,\infty)^{d}\mapsto[0,\infty)$ is a homogeneous function of order $1$ named stable-tail dependence function (Beirlant et al., 2004, chap. 8). This latter function fully characterizes the extremal dependence structure; however, since it is defined on $0,\infty)^{d}$ and takes values in $0,\infty)$, it is difficult to visualize directly. For this reason, it is common practice to work instead with its restriction to the set ${\mathcal{R}}:=\{{\bm{t}}\in[0,1]^{d-1}:\lVert{\bm{t}}\rVert_{1}\leqslant 1\}$, known as the Pickands dependence function (e.g., Falk et al., 2010, Ch. 4), which is more amenable to graphical representation. While no exhaustive characterization of the Pickands dependence function is available, it must at least satisfy the following convexity and boundary conditions
@@
In the bivariate case, these are necessary and sufficient to define a valid Pickands dependence function (for details, see Beirlant et al., 2004, chap. 8). However, there is an alternative way to represent extreme dependence. Indeed, a formal definition of the Pickands dependence function is
@@
where $H$ is a probability measure on the $d$-dimensional unit simplex $\mathcal{S}_{d}:=\{{\bm{w}}\geqslant{\bf 0}:\|{\bm{w}}\|_{1}=1\}$}, named angular measure (e.g., Falk et al., 2010, Ch. 4). Thus, the level of dependence can be conveyed through $H$; however, the same issue arises again: which specific measure should one consider? A valid angular measure is any probability measure satisfying the mean constraint
@@
The result in (2.1) can be interpreted as follows. Given a sample $({\bm{X}}_{1},\ldots,{\bm{X}}_{N})$ with joint distribution $F\in\mathcal{D}(G)$, for each of the $d$ variables we compute $k$ maxima are computed over blocks of size $n$, such that $N=n\cdot k$, where the $i$th maximum is $M_{i,j}^{(n)}=\max(X_{(i-1)n+1,j},\ldots,X_{in,j})$, $i\in\{1,\ldots,k\}$, $j=1,\ldots,d$. We bind the maxima in the vector of componentwise maxima ${\bm{M}}_{i}^{(n)}=(M_{i,1}^{(n)},\ldots,M_{i,d}^{(n)})^{\top}$. Now, setting ${\bm{\mu}}={\bm{b}}_{n}$, ${\bm{\sigma}}={\bm{a}}_{n}$, ${\bm{y}}={\bm{a}}_{n}{\bm{x}}+{\bm{b}}_{n}$, then thanks to (2.1), the joint distribution of vector of componentwise maxima can be approximated, for large enough $n$, as $F^{n}({\bm{y}})\approx G_{{\bm{\theta}}}({\bm{y}}\mid D),$ where ${\bm{\theta}}=({\bm{\mu}},{\bm{\sigma}},{\bm{\gamma}})^{\top}$, By construction, the distribution of the univariate maximum $M_{i,j}^{(n)}$ can be approximated by
$F^{n}(y_{j})\approx G_{\theta_{j}}(y_{j})=\exp\left\{-\left(1+\gamma_{j}\frac{y_{j}-\mu_{j}}{\sigma_{j}}\right)^{-\frac{1}{\gamma_{j}}}_{+}\right\},\qquad j\in\{1,\ldots,d\},$ (2.6)
@@
Recall that the joint distribution function $F$ can be represented in terms of a copula function, $C$, defined on $0,1]^{d}$ and its marginal distributions $F_{i},i=1,\ldots,d$ (assumed to be continuous) through $F({\bm{x}})=C\{F_{1}(x_{1}),\ldots,F_{d}(x_{d})\}$ (see, e.g., Joe, 2014, for details). The limiting result in formula (2.8) implies the following important approximation: let $p_{j}=x_{j}/n$, for $j=1,\ldots,d$, and $Q_{j}(p_{j})$ be the (1p1-p) quantile of $F_{j}$; then, by homogeneity of the stable tail dependence function, the probability that at least one among the $d$ variables $X_{j}$ exceeds a high quantile of its own distribution, can be approximated for large $n$ by
${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{or}\cdots\text{or}\,X_{d}>Q_{d}(p_{d})\}\approx L(p_{1},\ldots,p_{d}).$ (2.9)
@@
${\mathbb{P}}\{X_{1}>Q_{1}(p_{1})\,\text{and}\cdots\text{and}\,X_{d}>Q_{d}(p_{d})\}\approx R(p_{1},\ldots,p_{d}).$ (2.10)
@@
Finally, an important question is, how to determine an extreme region where two variables, given a small joint probability $p$ of falling in it? We summarize the approach discussed by Cai et al. (2011), Einmahl et al. (2013) and He and Einmahl (2017) as follows with a focus on the bivariate case explored by Einmahl et al. (2013). Let $f$ be the density on ${\mathbb{R}}_{+}^{2}$ of a bivariate distribution $F$ satisfying $F\in\mathcal{D}(G)$, with $\gamma_{1},\gamma_{2}>0$. A possible definition of quantile regions is given by the level sets of $f$, namely for $\alpha>0$
@@
Assume that for some $\mathcal{M}>0$, $f$ is bounded away from zero on $(0,\mathcal{M}]^{2}$ and, outside this region, $f$ is decreasing in each coordinate and that the exponent measure admits a density $g$ satisfying
@@
where $k=k_{N}$ with $k\to\infty$ as $N\to\infty$ and $k=o(N)$, which approximate the target extreme quantile region $\mathcal{Q}_{N}$. In particular, $\mu_{j}=Q_{j}(k/N)$ and $\sigma_{j}=a_{j}(N/k)$ with $j=1,2$, where $a_{j}(\cdot)$ is a suitable positive scale function (de Haan and Ferreira, 2006, Ch. 1–2). Then, the set $\widetilde{\mathcal{Q}}_{N}$ is a good approximation of $\mathcal{Q}_{N}$ in the sense that ${\mathbb{P}}(\widetilde{\mathcal{Q}}_{N})\approx p$ and ${\mathbb{P}}(\mathcal{Q}_{N}\Delta\widetilde{\mathcal{Q}}_{N})\to 0$ as $N\to\infty$ (see Einmahl et al., 2013), where $A\Delta B=A\setminus B\cup B\setminus A$.